sepidks tie Find lhe indicated zsore. The graph depicts the etandard noral dietr bution with mearn 0 and standard deviatio 1 The indicated z 9**re is Round to two decimal placas as neaded
Answer #1

Here,

Z ~ Normal(0,1)

To find z0 such that

P(Z >刎ー0.1230

Rightarrow 1 - P(Zleq z_0) = 0.1230

(Z-20) = 1-0. 1230 0.877

Φ(so) = 0.877

20-Φ-1 (0.877)

20 1.160 12(ans)